# Spacetime Measurements with the Photon Ring

### v2

@@ III Variation of Critical Parameters with Observer Inclination and Non-Kerr Spacetimes
IV Summary & Conclusions
Horizon-scale images of BHs have opened a new frontier for probing strong-field gravity and revealing key properties of astrophysical BHs, such as their spin. A particularly exciting prospect for future BH imaging, with higher angular resolution and greater flux sensitivity, is the direct detection of the BH photon ring. Due to strong gravitational lensing, a single point source can produce multiple images on the observer’s screen. The primary image is the first to arrive on the screen, while “higher-order” images – collectively known as photon ring for an extended optically thin source of emission – emerge later. These increasingly higher-order images are formed by photons that execute multiple (half-)orbits around the BH, carrying unique signatures of the near-horizon spacetime geometry.
@@ IV Summary & Conclusions
Thus, $\langle\bar{\gamma}\rangle_{\psi}\leq 0$, $\langle\bar{\delta}\rangle_{\psi}\geq 0$, and $\langle\bar{\tau}\rangle_{\psi}\leq 0$ imply that photon subrings in these non-Kerr spacetimes are wider, brighter, more rotated, and advanced compared to those in the Schwarzschild spacetime.
Quantitatively, for Kerr BHs, the fractional deviations across the full parameter space of spin and inclination are: $\sim 20\%$ in $\langle\bar{\gamma}\rangle_{\psi}$, $\sim 10\%$ in $\langle\bar{\tau}\rangle_{\psi}$, and $\sim 60\%$ in $\langle\bar{\delta}\rangle_{\psi}$ – significantly larger than the $8\%$ variation in the shadow diameter from their corresponding values for Schwarzschild BH.
@@
Each critical parameter offers an independent measurement of a distinct spacetime property near the BH photon shell. The demagnification exponent is sensitive to spacetime curvature, the time delay reflects the gravitational potential governing time dilation, and the rotation parameter provides insight into spacetime frame-dragging. Therefore, joint measurements of these critical parameters can help deduce key BH properties.
While critical parameters provide a powerful framework for characterizing photon ring and BH features, potential degeneracies between different spacetime solutions may limit their ability to uniquely identify the underlying spacetime geometry and deviations (e.g., charge) from the Kerr geometry.
@@ Acknowledgements.
RKW’s research is supported by the Fulbright-Nehru Postdoctoral Research Fellowship (Award No. 2847/FNPDR/2022) from the United States-India Educational Foundation. PK acknowledges support from grants from the Gordon and Betty Moore Foundation (GBMF-8273) and the John Templeton Foundation (#62286) to the BH Initiative at Harvard University. The Perimeter Institute for Theoretical Physics partially supported this work. Funding for research at the institute is provided by the Department of Innovation, Science and Economic Development Canada, and the Ministry of Economic Development, Job Creation and Trade of Ontario, both of which are branches of the Government of Canada. DOC acknowledges financial support from the Brinson Foundation, the Gordon and Betty Moore Foundation (GBMF-10423), and the National Science Foundation (AST-2307887, AST-1935980, and AST-2034306). This work was supported by the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (Grant #62286) and the Gordon and Betty Moore Foundation (Grant GBMF-8273) - although the opinions expressed in this work are those of the author(s) and do not necessarily reflect the views of these Foundations.
@@ Appendix A Representative Measures of The critical Parameters
Figure 6: Variation of fractional deviation in critical parameters for the Kerr BH (panels [a.], [c.], [e.], [g.]) and Kerr-Newman BH with $Q=0.4$ (panels [b.], [d.], [f.], [h.]). All deviations are shown with respect to their Schwarzschild values over varying image plane polar angle $\psi$; the demagnification exponent, $\bar{\gamma}$, (panels [a.] and [b.]), the time delay between successive images, $\bar{\tau}$, ([c.] and [d.]), the rotation parameter, $\bar{\delta}$, ([e.] and [f.]), and the phase wrapped rotation parameter, $\bar{\delta}_{2\pi}$, ([g.] and [h.]). We show deviation curves for varying spin values varying from $0.01\%$ to $99\%$ of extremality for spacetimes in varying colors, and for observer inclination angles of an M87-like system of $17^{\circ}$ (solid), and for an extreme inclination case of $89^{\circ}$ (dashed). The insets of [d.] and [f.] displayed a zoomed in version of their encompassing panels.
@@
While $\bar{\gamma}<0$ for all values of $a$ and $\mathscr{i}$, at high inclinations, $\bar{\gamma}$ exhibits non-monotonic variation between $\psi=0$ and $\psi=\pi$, reaching a minimum at an intermediate value. This minimum in $\bar{\gamma}$ corresponds to the screen angle where the photon ring appears widest and brightest. Both $\bar{\delta}\geq 0$ and $\bar{\tau}\geq 0$ further increase with the screen angle $\psi$. This monotonically increasing time delay with $\psi$ is an indicator that the distinct part of a given order photon ring takes distinct time to appear. This temporal variation arises because prograde photons $r_{\rm p}\leq r_{\rm p}^{0}$ forming a part of photon ring takes longer to reach the observer’ screen due to time dilation, compared to their retrograde counterparts. Consequently, for a counter-clockwise spinning BH, the left part of the photon ring ($\pi/2\leq\psi\leq 3\pi/2$) on the image plane appears more wider and brighter, more rotated, and delayed than its right part.

### v3

@@ III Variation of Critical Parameters with Observer Inclination and Non-Kerr Spacetimes
IV Summary And Conclusions
Horizon-scale images of BHs have opened a new frontier for probing strong-field gravity and revealing key properties of astrophysical BHs, such as their spin. A particularly exciting prospect for future BH imaging, with better angular resolution and greater flux sensitivity, is the direct detection of the BH photon ring. Due to strong gravitational lensing, a single point source can produce multiple images on the observer’s screen. The primary image is the first to arrive on the screen, while “higher-order” images—collectively known as photon ring for an extended optically thin source of emission—emerge later. These increasingly higher-order images are formed by photons that execute multiple (half-)orbits around the BH, carrying unique signatures of the near-horizon spacetime geometry.
The novelty of this work lies in extending the analyses of Refs. [16] and [52]. While Ref. [16] derives critical parameters exclusively for Kerr BHs and Ref. [52] examines only the shadow size and Lyapunov exponent for a polar observer, this study explores how strong-field deviations from the Kerr spacetime affect all critical parameters across varying inclinations.
@@ IV Summary And Conclusions
Thus, $\langle\bar{\gamma}\rangle_{\psi}\leq 0$, $\langle\bar{\delta}\rangle_{\psi}\geq 0$, and $\langle\bar{\tau}\rangle_{\psi}\leq 0$ imply that photon subrings in these non-Kerr spacetimes are wider, brighter, more rotated, and advanced in time compared to those in the Schwarzschild spacetime.
Quantitatively, for Kerr BHs, the fractional deviations across the full parameter space of spin and inclination are: $\sim 20\%$ in $\langle\bar{\gamma}\rangle_{\psi}$, $\sim 10\%$ in $\langle\bar{\tau}\rangle_{\psi}$, and $\sim 60\%$ in $\langle\bar{\delta}\rangle_{\psi}$—significantly larger than the $8\%$ variation in the shadow radius from their corresponding values for Schwarzschild BH.
@@
Building on this, we note that each critical parameter offers an independent measurement of a distinct spacetime property near the BH photon shell. The demagnification exponent is sensitive to spacetime curvature, the time delay reflects the gravitational potential governing time dilation, and the rotation parameter provides insight into spacetime frame-dragging. Therefore, joint measurements of these critical parameters can help deduce key BH properties.
Encouragingly, with the growing capabilities of interferometric arrays, such as ngEHT [23] and BHEX [24], the precision required to detect $n=1$ photon ring is now within reach, bringing us closer to addressing fundamental questions about the nature of BHs and their potential deviations from the Kerr spacetimes.
Nonetheless, while critical parameters provide a powerful framework for characterizing photon ring and BH features, potential degeneracies between different spacetime solutions may limit their ability to uniquely identify the underlying spacetime geometry and deviations (e.g., charge) from the Kerr geometry.
Motivated by these findings, this work opens new research directions: (i) assessing the accuracy of critical parameters in defining the scaling relations between $n=0$ and $n=1$ rings, (ii) how precisely the calibrated critical parameters, that define the scaling relations between $n=0$ and $n=1$ rings, determine the BH parameters. Our future research is exploring these questions.
@@ Acknowledgements.
R.K.W.’s research is supported by the Fulbright-Nehru Postdoctoral Research Fellowship (Award No. 2847/FNPDR/2022) from the United States-India Educational Foundation. P.K. acknowledges support from grants from the Gordon and Betty Moore Foundation (No. GBMF-8273) and the John Templeton Foundation (No. 62286) to the Black Hole Initiative at Harvard University. D.O.C acknowledges financial support from the Brinson Foundation, the Gordon and Betty Moore Foundation (No. GBMF-10423), and the National Science Foundation (No. AST-2307887, No. AST-1935980, and No. AST-2034306). K.S. acknowledges support from the Perimeter Institute for Theoretical Physics. Funding for research at the institute is provided by the Department of Innovation, Science and Economic Development Canada, and the Ministry of Economic Development, Job Creation and Trade of Ontario, both of which are branches of the Government of Canada. The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these foundations.
@@ Appendix A Representative Measures of The critical Parameters
Figure 6: Variation of fractional deviation in critical parameters for the Kerr BH [panels (a), (c), (e), (g)] and Kerr-Newman BH with $Q=0.4$ [panels (b), (d), (f), (h)]. All deviations are shown with respect to their Schwarzschild values over varying image plane polar angle $\psi$; the demagnification exponent, $\bar{\gamma}$, [panels (a) and (b)], the time delay between successive images, $\bar{\tau}$, [(c) and (d)], the rotation parameter, $\bar{\delta}$, [(e) and (f)], and the phase wrapped rotation parameter, $\bar{\delta}_{2\pi}$, [(g) and (h)]. We show deviation curves for varying spin values varying from $0.01\%$ to $99\%$ of extremality for spacetimes in varying colors, and for observer inclination angles of an M87-like system of $17^{\circ}$ (solid), and for an extreme inclination case of $89^{\circ}$ (dashed). The insets of (d) and (f) displayed a zoomed in version of their encompassing panels.
@@
While $\bar{\gamma}<0$ for all values of $a$ and $\mathscr{i}$, at high inclinations, $\bar{\gamma}$ exhibits nonmonotonic variation between $\psi=0$ and $\psi=\pi$, reaching a minimum at an intermediate value. This minimum in $\bar{\gamma}$ corresponds to the screen angle where the photon ring appears widest and brightest. Both $\bar{\delta}\geq 0$ and $\bar{\tau}\geq 0$ further increase with the screen angle $\psi$. This monotonically increasing time delay with $\psi$ is an indicator that the distinct part of a given order photon ring takes distinct time to appear. This temporal variation arises because prograde photons $r_{\rm p}\leq r_{\rm p}^{0}$ forming a part of photon ring takes longer to reach the observer’ screen due to time dilation, compared to their retrograde counterparts. Consequently, for a counterclockwise spinning BH, the left part of the photon ring ($\pi/2\leq\psi\leq 3\pi/2$) on the image plane appears more wider and brighter, more rotated, and delayed than its right part.